OpenStudy (anonymous):

Please help me check my work. I needed to find the tangent plane and normal line of the function f(x,y) = x^3*y at point (1,2,f(1,2)).

OpenStudy (anonymous):

I wrote for the plane z = f(1,2) + a (x - 1) + b (y - 2) where a and b are the partial derivatives of f(x,y) at the point (1,2) f(1,2) = 2 a = 3x^2 y at the point (1,2) = 6 b = x^3 at the point (1,2) = 1 z = 2 + 6 (x - 1) + (y - 2) Then I chose two points on the plane: (1,2,2) and (2,-4,2) and subtracted the second from the first getting: (-1,6,0), so (-1,6) is parallel to the plane, and the normal line goes through (1,2,2) and is normal to it. r: (x,y,z) = (1,2,2) + lambda (6,1,0)

OpenStudy (turingtest):

I would just rewrite the surface as f(x,y)=z then create a new surface function in R3 g(x,y,z)=x^3y-z find grad(g) at x=1, y=2 (which from the function we know is also at z=2) that is your normal vector n

OpenStudy (turingtest):

looks like you got the normal vector to be n=(6,1,0) but I got n=(6,1,-1)

OpenStudy (turingtest):

\[g(x,y,z)=x^3y-z\]\[\nabla g=\langle3x^2y,x^3,-1\rangle\]\[\nabla g(1,2,2)=\langle6,1,-1\rangle=\vec n\]the given point on the surface is\[\vec P_0=\langle1,2,2\rangle\]and let\[\vec P=\langle x,y,z\rangle\]now you can use\[\vec n\cdot\vec P_0=\vec n\cdot\vec P\]to get the equation of the plane

OpenStudy (turingtest):

so your plane is right, but I think your normal line is wrong use\[\vec r(t)=\vec P_0+t\vec n\]to find the equation of the line

OpenStudy (anonymous):

Thanks.

OpenStudy (turingtest):

welcome :)